HIGH PRECISION MEASUREMENT OF C-O STRETCHING BAND OF CH_{3}OH USING CO_{2} LASER SIDE-BANDS

\underline{Z. D. SUN}, F. MATSUSHIMA, S. TSUNEKAWA, and K. TAKAGI, Department of Physics, Toyama University, Toyama 930-8555, Japan..

Microwave sidebands of CO_{2} laser lines have been used as an infrared power source to measure the sub-Doppler vibrationrotation frequencies of CH_{3}OH in the C-O stretching band precisely. More than one hundred frequencies of CH_{3}OH have been measured. Over thirty groups of unresolved lines in the Fourier-transform spectrum have been resolved. Combined with our previous infrared-microwave double resonance results^{a} and the ground state (υ_{co} = 0) term values given in literature^{b}, many term valves of CH_{3}OH in the υ_{co} = 1 stretching states have been refined. For A species in υ_{co} = 1, 22 asymmetry splittings for various K in υ_{t} = 0 and 1 have been observed for the first time. The present results provide precision data to refine the molecular constants in the first excited C-O stretching state of CH_{3}OH.

aZ. D. Sun, S. Ogura, F. Matsushima, S. Tsunekawa and K. Takagi, paper RA08 at the “53^{rd} Ohio State University International Symposium on Molecular Spectroscopy”, Columbus, Ohio (1998).

bG. Moruzzi, B. P. Winnewisser, M. Winnewisser, I. Mukhopadhyay, and F. Strumia, Microwave, infrared and laser transitions of methanol, (CRC Press, Inc. 1995).